"Mikey takes 4 hours to complete a job. Marko takes 5 hours to complete the same job. How much time will it take to complete the job if they work together?"
I know this is a fractional word problem, however I don't know how to set it up properly and solve. Can I get help with solving?
Thank You.

Answer by richwmiller(17219)

x/4+x/5=1
5x+4x=20
9x=20
x=20/9
x= 2 2/9 hours